• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 08:01:49 on April 30, 2007

    Entry number 200608

    keyword=Owl Shift Summary

    Owl Shift for Monday April 30th, 2007


    Leader:   Bryan Moffit
    Target:   Elena Khrosinkova
    Beam Energy = 4.501 GeV
    Left arm: Angle = 21.546° p0 = 3.728 GeV/c
    Right arm: Angle = 68.328° p0 = 0.727 GeV/c

    Summary

    Smooth production data taking.

    Significant Events

    • 00:00 Began shift in the middle of Run 2999. 100µA.
    • 00:15 MCC confirms that the target OTR was removed around 23:50.
    • 04:56 Move to Al dummy target.
    • 05:05 Request 40µA. MCC had to flip FFB off and on so that the BPMs could do a gain search at the lower current.
    • 06:30 Move to 4He target.
    • 06:38 Request 100µA.
    • 08:00 Pass Run 3014 to Day Shift.

    Data Run Summary

    • 00:00 Run 2999 - Production Run (4He)
    • 00:12 Run 3000 - Production Run (4He)
    • 00:38 Run 3001 - Production Run (4He)
    • 01:06 Run 3002 - Production Run (4He)
    • 01:33 Run 3003 - Production Run (4He)
    • 01:59 Run 3004 - Production Run (4He)
    • 02:30 Run 3005 - Production Run (4He)
    • 02:58 Run 3006 - Production Run (4He)
    • 03:27 Run 3007 - Production Run (4He)
    • 04:00 Run 3008 - Production Run (4He)
    • 04:31 Run 3009 - Production Run (4He)
    • 05:17 Run 3010 - Background Run (Al dummy). Stopped early to change prescale factors.
    • 05:33 Run 3011 - Background Run (Al dummy).
    • 06:47 Run 3012 - Production Run (4He)
    • 07:19 Run 3013 - Production Run (4He)
    • 07:45 Run 3014 - Production Run (4He)


    You went over my helmet?